# InterCLIP-MEP: Interactive CLIP and Memory-Enhanced Predictor for Multi-modal Sarcasm Detection

### v4

@@ Introduction
21:     $\textrm{logits}\leftarrow[\sum_{k=0}^{\mathcal{I}[0]}(\hat{h}^{f}_{i}\mathcal{M}[0]^{T})_{k},\sum_{k=0}^{\mathcal{I}[1]}(\hat{h}^{f}_{i}\mathcal{M}[1]^{T})_{k}]$
@@
The detailed computational process of MEP is provided in Algorithm 1, where $N_{\textrm{test}}$ denotes the number of test samples. MEP uses the trained InterCLIP to extract fused features of the samples. It utilizes the classification module $\mathcal{F}_{c}$ to assign a pseudo-label $\ell_{\textrm{pse}_{i}}$ to each sample $\mathcal{P}_{i}$ and the projection module $\mathcal{F}_{p}$ to obtain the sample’s projected feature $\hat{h}^{f}_{i}$. To store valuable projected features of test samples as historical knowledge, MEP maintains a dynamic fixed-length dual-channel memory $\mathcal{M}\in\mathcal{R}^{2\times L\times d_{f}}$, where $L$ is the memory length per channel. The first channel stores projected features of non-sarcastic samples, while the second channel stores those of sarcastic samples. Based on the pseudo-label $\ell_{\textrm{pse}_{i}}$, the appropriate memory channel $\mathcal{M}[\ell_{\textrm{pse}_{i}}]$ is selected for updating. If the selected channel has available space, the sample’s projected features are added directly, and the prediction entropy is recorded. If the memory is full, the prediction entropy of all samples in the memory is compared with that of the current sample, replacing the samples with the highest entropy as necessary. Finally, the current sample’s projected features are combined with the historical features stored in both memory channels $\mathcal{M}$ using cosine similarity to yield the final prediction.
All Sarcastic Non-sarcastic
Train 19,812 9,572 10,240
Validation 2,410 1,042 1,368
Test 2,409 1,037 1,372
Table 1: Statistics of the dataset from the MMSD2.0 benchmark (Qin et al. 2023).
Modality Method Acc. (%) F1 (%) P (%) R (%)
Text TextCNN* (Kim 2014) 71.61 69.52 64.62 75.22
Bi-LSTM* (Graves and Schmidhuber 2005) 72.48 68.05 68.02 68.08
SMSD* (Xiong et al. 2019) 73.56 69.97 68.45 71.55
RoBERTa* (Liu et al. 2019) 79.66 76.21 76.74 75.70
Image ResNet* (He et al. 2016) 65.50 57.58 61.17 54.39
ViT* (Dosovitskiy et al. 2021) 72.02 69.72 65.26 74.83
Text-Image HFM* (Cai, Cai, and Wan 2019) 70.57 66.88 64.84 69.05
Att-BERT* (Pan et al. 2020) 80.03 77.04 76.28 77.82
CMGCN* (Liang et al. 2022) 79.83 76.90 75.82 78.01
HKE* (Liu, Wang, and Li 2022) 76.50 72.25 73.48 71.07
DIP (Wen, Jia, and Yang 2023) 80.59 78.23 75.52 81.14
DynRT (Tian et al. 2023) 70.37 68.55 63.02 75.15
Multi-view CLIP* (Qin et al. 2023) 85.64 84.10 80.33 88.24
G2SAM (Wei et al. 2024) 79.43 78.07 72.04 85.20
InterCLIP-MEP w/o Inter ($L=1024$) 86.05 84.81 79.83 90.45
InterCLIP-MEP w/ TW ($L=128$) 85.51 84.26 79.15 90.07
InterCLIP-MEP w/ V2T ($L=640$) 86.26 85.00 80.17 90.45
InterCLIP-MEP w/ T2V ($L=1024$) 86.72 85.61 80.20 91.80
@@
Table 2: Main results. We use * to indicate that the results are taken from Qin et al. (2023). $\dagger$ indicates the results obtained by running this method on the MMSD2.0 benchmark (Qin et al. 2023). Underlined values represent the best baseline for comparison. Bold values indicate those that surpass the baseline.
@@
Table 3: Ablation study of InterCLIP-MEP. Baseline represents the results of the corresponding method without any ablation.
Experiment
Experimental Settings
Datasets.
We use the dataset from the MMSD2.0 benchmark (Qin et al. 2023), which is an improved version of the dataset originally introduced by Cai, Cai, and Wan (2019). We present the statistics of the dataset in Table 1.
Metrics.
Following Liu, Wang, and Li (2022), we use accuracy (Acc.), precision (P), recall (R), and F1-score (F1) as metrics to evaluate the performance of our model.
@@
We compare the effectiveness of the InterCLIP-MEP framework against several unimodal and multi-modal methods. For text modality methods, we compare with TextCNN (Kim 2014), Bi-LSTM (Graves and Schmidhuber 2005), SMSD (Xiong et al. 2019), and RoBERTa (Liu et al. 2019). For image modality methods, we compare with ResNet (He et al. 2016) and ViT (Dosovitskiy et al. 2021). For state-of-the-art multi-modal methods, we compare with the following:
  • HFM (Cai, Cai, and Wan 2019): A hierarchical fusion model integrating text and image features.
  • Att-BERT (Pan et al. 2020): A model using an attention mechanism to capture sarcasm incongruity.
  • CMGCN (Liang et al. 2022): A method that constructs a cross-modal graph and employs a graph convolutional network.
  • HKE (Liu, Wang, and Li 2022): A hierarchical framework utilizing multi-head cross-attention and graph neural networks.
  • DIP (Wen, Jia, and Yang 2023): A dual incongruity perceiving network that extracts sarcastic information from factual and affective levels.
  • DynRT (Tian et al. 2023): A dynamic routing transformer network for multi-modal sarcasm detection.
  • Multi-view CLIP (Qin et al. 2023): A method adapted to CLIP for capturing multi-view sarcasm cues.
  • G2SAM (Wei et al. 2024): A method using global graph-based semantic awareness and leveraging historical knowledge from training samples for multi-modal sarcasm detection.
Main Results
@@
To validate the effectiveness of our InterCLIP-MEP framework, we conduct experiments using the original CLIP as the backbone instead of InterCLIP, referred to as InterCLIP-MEP w/o Inter. We compare this configuration with three interaction modes of InterCLIP: InterCLIP-MEP w/ V2T, InterCLIP-MEP w/ T2V, and InterCLIP-MEP w/ TW.
@@
For each experiment, we condition only the top four layers of the self-attention modules, with the projection dimension $d_{f}$ set to 1024. We set the rank $r$ of LoRA to 8, fine-tuning the self-attention module weight matrices $\mathbf{W}$, specifically $W_{k}$, $W_{v}$, and $W_{o}$. For the memory size $L$ of the MEP, we select the optimal size from $\mathbf{L}=\{128,256,384,512,640,768,896,1024,1152,1280\}$.
@@
The main results are shown in Table 2. Our framework consistently outperforms or matches the performance of state-of-the-art methods, whether using InterCLIP or the original CLIP as the backbone, as shown in Table 2. This demonstrates the effectiveness of our training strategy and MEP. Methods InterCLIP-MEP w/ V2T and InterCLIP-MEP w/ T2V outperform InterCLIP-MEP w/o Inter, indicating that InterCLIP more effectively captures text-image interactions. Additionally, InterCLIP-MEP w/ T2V outperforms InterCLIP-MEP w/ V2T, likely due to the complexity of the visual space, which poses challenges for the projection layer in mapping visual representations into the text encoder space effectively. We observe that InterCLIP-MEP w/ TW performs worse than InterCLIP-MEP w/o Inter, possibly because embedding representations within both encoders increases learning difficulty compared to unidirectional methods. In summary, InterCLIP with T2V interaction, combined with our training strategy and MEP yields the most promising results.
Ablation Study
@@
To further validate the effectiveness of InterCLIP-MEP, we remove the projection module $\mathcal{F}_{p}$ and train only the classification module $\mathcal{F}_{c}$ for prediction, denoted as w/o Proj. To test the necessity of using LoRA (Hu et al. 2022) for fine-tuning, we keep the rest of InterCLIP-MEP unchanged and freeze all self-attention weight matrices of InterCLIP, denoted as w/o LoRA, always selecting the optimal memory size $L$ for the MEP during inference. To evaluate the effectiveness of the MEP, we train both $\mathcal{F}_{p}$ and $\mathcal{F}_{c}$ but use only $\mathcal{F}_{c}$ during inference, denoted as w/o MEP.
@@
Table 3 reports all results. All variants show performance declines compared to the baseline, demonstrating the importance of each module in the InterCLIP-MEP framework. For methods InterCLIP-MEP w/ TW and InterCLIP-MEP w/o Inter, the w/o MEP variant performed worse than the w/o Proj variant. However, for methods InterCLIP-MEP w/ T2V and InterCLIP-MEP w/ V2T, the w/o MEP variant performs better than the w/o Proj variant. This suggests that backbones with strong image-text interaction capabilities benefit from training the classification module $\mathcal{F}_{c}$ along with the projection module $\mathcal{F}_{p}$, even without using MEP during inference. We also find that not using LoRA to fine-tune the self-attention modules results in significant performance loss, indicating that the original CLIP’s vanilla space is not directly suitable for the sarcasm detection task.
Hyperparameter Study
In this section, we further investigate the method using Interactive-CLIP with T2V interaction as the backbone. Keeping the other hyperparameters constant, we condition different top-$n$ layers of the self-attention modules. We also study the impact of different projection dimensions $d_{f}$, different LoRA ranks $r$, and different memory sizes $L$ on the InterCLIP-MEP w/ T2V method. We present all results in Figure 4.
Figure 4: Hyperparameter study curves for InterCLIP-MEP w/ T2V. Panel (d) compares results with those from using only the classification module $\mathcal{F}_{c}$ for prediction.
Figure 4(a) shows that conditioning the top four self-attention layers yields the best results. From Figure 4(b), a rank of 8 is optimal. Figure 4(c) indicates the projection dimension is best at 256 or 1024. Figure 4(d) reveals that a memory size of 640 in MEP outperforms others, confirming the value of historical sample knowledge.
Figure 5: Case study of InterCLIP-MEP.
Figure 6: Visual comparison between InterCLIP and CLIP.
Case Study and Visualization

### v5

@@ 3.1 Interactive CLIP
21:     $\textrm{logits}\leftarrow\left[\sum_{k=0}^{\mathcal{I}[0]}(\hat{h}^{f}_{i}\mathcal{M}[0]^{T})_{k},\sum_{k=0}^{\mathcal{I}[1]}(\hat{h}^{f}_{i}\mathcal{M}[1]^{T})_{k}\right]$
@@
3.2 Memory-Enhanced Predictor
As depicted in Figure 2 (right), we present the Memory-Enhanced Predictor (MEP) that builds upon the learned InterCLIP, along with the classification module and the projection module, leveraging the valuable historical knowledge of test samples to enhance multi-modal sarcasm detection.
@@ 3.2 Memory-Enhanced Predictor
The detailed computational process of MEP is provided in Algorithm 1, where $N_{\textrm{test}}$ denotes the number of test samples. MEP uses the trained InterCLIP to extract fused features of the samples. It utilizes the classification module $\mathcal{F}_{c}$ to assign a pseudo-label $\ell_{\textrm{pse}_{i}}$ to each sample $\mathcal{P}_{i}$ and the projection module $\mathcal{F}_{p}$ to obtain the sample’s projected feature $\hat{h}^{f}_{i}$. To store valuable projected features of test samples as historical knowledge, MEP maintains a dynamic fixed-length dual-channel memory $\mathcal{M}\in\mathcal{R}^{2\times L\times d_{f}}$, where $L$ is the memory length per channel. The first channel stores projected features of non-sarcastic samples, while the second channel stores those of sarcastic samples. Based on the pseudo-label $\ell_{\textrm{pse}_{i}}$, the appropriate memory channel $\mathcal{M}[\ell_{\textrm{pse}_{i}}]$ is selected for updating. If the selected channel has available space, the sample’s projected features are added directly, and the prediction entropy is recorded. If the memory is full, the prediction entropy of all samples in the memory is compared with that of the current sample. Samples with the highest entropy are replaced, ensuring the retained samples have lower entropy. Finally, the current sample’s projected feature is combined with the historical features stored in both memory channels $\mathcal{M}$ using cosine similarity to yield the final prediction.
4 Experiment
MMSD/MMSD2.0 Sarcastic Non-sarcastic All
Train 8,642/9,576 11,174/10,240 19,816/19,816
Validation 959/1,042 1,451/1,368 2,410/2,410
Test 959/1,037 1,450/1,372 2,409/2,409
Table 1: Statistics of MMSD and MMSD2.0.
Method MMSD2.0 MMSD
Acc. (%) F1 (%) P (%) R (%) Acc. (%) F1 (%) P (%) R (%)
Text
TextCNN Kim (2014) 71.61* 69.52* 64.62* 75.22* 80.03* 75.32* 74.29* 76.39*
Bi-LSTM Zhou et al. (2016) 72.48* 68.05* 68.02* 68.08* 81.90* 77.53* 76.66* 78.42*
SMSD Xiong et al. (2019) 73.56* 69.97* 68.45* 71.55* 80.90* 75.82* 76.46* 75.18*
RoBERTa Liu et al. (2019) 79.66* 76.21* 76.74* 75.70* 93.97* 92.45* 90.39* 94.59*
Image
ResNet He et al. (2016) 65.50* 57.58* 61.17* 54.39* 64.76* 61.53* 54.41* 70.80*
ViT Dosovitskiy et al. (2021) 72.02* 69.72* 65.26* 74.83* 67.83* 63.40* 57.93* 70.07*
Text-Image
HFM Cai et al. (2019) 70.57* 66.88* 64.84* 69.05* 83.44* 80.18* 76.57* 84.15*
Att-BERT Pan et al. (2020) 80.03* 77.04* 76.28* 77.82* 86.05* 82.92* 80.87* 85.08*
CMGCN Liang et al. (2022) 79.83* 76.90* 75.82* 78.01* 86.54* 84.09* -* -*
HKE Liu et al. (2022) 76.50* 72.25* 73.48* 71.07* 87.36* 72.25* 81.84* 86.48*
DIP Wen et al. (2023) 80.59 78.23 75.52 81.14 89.59 87.17 87.76 86.58
DynRT Tian et al. (2023) 70.37 68.55 63.02 75.15 93.59 91.93 90.30 93.62
Multi-view CLIP Qin et al. (2023) 85.64* 84.10* 80.33* 88.24* 88.33* 85.55* 82.66* 88.65*
G2SAM Wei et al. (2024) 79.43 78.07 72.04 85.20 90.48 88.48 87.95 89.02
InterCLIP-MEP (Ours)
w/o Inter ($L_{2}=1024,L_{1}=1280$) 86.05 84.81 79.83 90.45 88.75 86.31 83.73 89.05
w/ TW ($L_{2}=128,L_{1}=1152$) 85.51 84.26 79.15 90.07 88.54 86.32 82.25 90.82
w/ V2T ($L_{2}=640,L_{1}=1024$) 86.26 85.00 80.17 90.45 88.92 86.66 83.21 90.41
w/ T2V ($L_{2}=1024,L_{1}=1152$) 86.72 85.61 80.20 91.80 88.83 86.37 84.02 88.84
InterCLIP-MEP w/ RoBERTa (Ours)
w/o Inter ($L_{2}=640,L_{1}=128$) 77.21 75.55 70.20 81.77 93.94 92.54 90.77 94.37
w/ TW ($L_{2}=896,L_{1}=256$) 81.98 80.78 74.69 87.95 93.73 92.28 90.48 94.16
w/ V2T ($L_{2}=640,L_{1}=512$) 76.96 75.26 69.98 81.39 93.94 92.54 90.69 94.47
w/ T2V ($L_{2}=1024,L_{1}=384$) 82.81 81.55 75.81 88.24 93.73 92.28 90.56 94.06
@@ 4 Experiment
Table 2: Main results. We use * to indicate that the results are taken from Qin et al. (2023). - indicates that results are not reported. $\dagger$ indicates our reproduced results. Underlined values represent the best multi-modal baseline for comparison. Bold values indicate those that surpass the underlined baseline. $L_{2}$ for MMSD2.0 and $L_{1}$ for MMSD denote the optimal MEP memory sizes.
Method Accuracy (%) Trainable Parameters (M) Fitting Time / Epoch (s) Inference Time (s) GPU Memory Peak (GB)
Multi-view CLIP  85.64  165  488  51  15.59
DIP  80.59  196  OOM  OOM  OOM
G2SAM  79.43  116  90  13  18.32
DynRT  70.37  25  370  26  8.03
InterCLIP-MEP  86.72  8  55  6  6.14
Table 3: Efficiency comparison of different methods. To demonstrate the efficiency of InterCLIP-MEP, we selected several recent baselines for comparison. The analysis was conducted using the MMSD2.0 dataset on a single NVIDIA RTX 4090 GPU with a batch size of 128. In the table, Fitting Time / Epoch indicates the time required for each epoch during training and validation and OOM indicates Out of Memory, referring to GPU memory overflow.
@@
Table 4: Ablation study of InterCLIP-MEP, with Baseline denoting results without ablation.
4.1 Experimental Settings
Datasets and metrics.
Following Qin et al. (2023), we evaluate performance on MMSD Cai et al. (2019) and MMSD2.0 Qin et al. (2023) using accuracy (Acc.), precision (P), recall (R), and F1-score (F1) as metrics. We present the statistics of the two datasets in Table 1.
@@ 4.1 Experimental Settings
We compare the effectiveness of the InterCLIP-MEP framework against several unimodal and multi-modal methods. For text modality methods, we compare with TextCNN Kim (2014), Bi-LSTM Zhou et al. (2016), SMSD Xiong et al. (2019), and RoBERTa Liu et al. (2019). For image modality methods, we compare with ResNet He et al. (2016) and ViT Dosovitskiy et al. (2021). We compare with state-of-the-art multi-modal methods, including HFM Cai et al. (2019), Att-BERT Pan et al. (2020), CMGCN Liang et al. (2022), HKE Liu et al. (2022), DIP Wen et al. (2023), DynRT Tian et al. (2023), Multi-view CLIP Qin et al. (2023), and G2SAM Wei et al. (2024), which employ various techniques such as hierarchical fusion, graph neural networks, and dynamic routing for multi-modal sarcasm detection.
4.2 Main Results
@@
To validate the effectiveness of our InterCLIP-MEP framework, we conduct experiments using the original CLIP as the backbone instead of InterCLIP, referred to as w/o Inter. We compare this configuration with three interaction modes of InterCLIP: w/ V2T, w/ T2V, and w/ TW.
@@
For each experiment, we condition only the top four layers of the self-attention modules, with the projection dimension $d_{f}$ set to 1024. We set the LoRA rank $r$ to 8, fine-tuning the self-attention weight matrices $\mathbf{W}$, specifically $W_{k}$, $W_{v}$, and $W_{o}$. For the memory size $L$, we select the optimal size from a fixed set of candidate values $\mathbf{L}$  [^2] The fixed candidate values are {128, 256, 384, 512, 640, 768, 896, 1024, 1152, 1280}.. The main results are shown in Table 2.
Performance on MMSD2.0.
@@
For MMSD2.0, our framework consistently outperforms or matches the performance of state-of-the-art methods, whether using InterCLIP or the original CLIP as the backbone, as shown in Table 2 (InterCLIP-MEP). This demonstrates the effectiveness of our training strategy and MEP. Our results show that w/ V2T and w/ T2V outperform w/o Inter, demonstrating that InterCLIP captures text-image interactions more effectively. Furthermore, w/ T2V achieves superior performance compared to w/ V2T, likely due to the inherent complexity of the visual space, which presents challenges for the projection layer when mapping vision representations into the text encoder space. In contrast, w/ TW performs worse than other configurations, possibly because embedding representations within both encoders increases the learning difficulty. In summary, InterCLIP with T2V interaction, combined with our training strategy and MEP, delivers the most promising results. These findings underscore the robustness and adaptability of our framework, establishing it as a highly effective solution for capturing nuanced text-image interactions and addressing the complexities of multi-modal sarcasm detection.
Performance on MMSD.
For MMSD, the RoBERTa-based text modality baseline significantly outperforms other methods due to spurious cues in the text, enabling accurate predictions solely based on textual features Qin et al. (2023). Consequently, models like DynRT, G2SAM, and DIP, which utilize RoBERTa or BERT for text feature extraction, achieve high performance on MMSD but experience a significant drop on MMSD2.0. To further investigate, we conduct an additional experiment, InterCLIP-MEP w/ RoBERTa, replacing the original text encoder with RoBERTa. While this change led to state-of-the-art performance on MMSD, it disrupted InterCLIP’s modality alignment, causing a reasonable performance drop on MMSD2.0. This suggests that MMSD’s text data contains spurious cues that allow models to rely heavily on the text encoder, while MMSD2.0, having been cleaned, requires more robust multi-modal capabilities. We further find that the w/ V2T variant consistently outperforms others in both InterCLIP-MEP and InterCLIP-MEP w/ RoBERTa experiments, underscoring the model’s tendency to overly depend on text modality.
Efficiency Comparison.
Our training strategy demonstrates both remarkable effectiveness and outstanding efficiency. To validate this, we conducted a comprehensive comparative analysis against leading state-of-the-art methods, as detailed in Table 3. For instance, the Multi-view CLIP method Qin et al. (2023) employs a multi-layer Transformer encoder for feature fusion, which, while effective, introduces a significant number of trainable parameters. This results in slower training and inference speeds and greater memory consumption. Similarly, the DIP method Wen et al. (2023) caches historical samples during training, which hinders its ability to support large-batch training under limited resource conditions. In contrast, our method operates with a batch size of 128 while utilizing a significantly smaller number of trainable parameters, which translates to notably faster training and validation cycles. Furthermore, by incorporating minimal parameter modifications to adapt CLIP and utilizing simple yet effective linear layers for representation fusion, our approach achieves superior inference speeds and drastically reduced memory consumption. These results highlight the practicality of our framework, establishing it as a benchmark for both computational efficiency and performance in multi-modal sarcasm detection.
4.3 Analysis of InterCLIP-MEP
To robustly validate the effectiveness of InterCLIP-MEP, we conduct comprehensive ablation studies and case studies on the more reliable MMSD2.0 benchmark, offering deeper insights into its design and performance. In addition, we include visualization analyses to provide an intuitive understanding of how the framework processes multi-modal sarcasm cues.
Ablation study.
@@ 4.3 Analysis of InterCLIP-MEP
We remove the projection module $\mathcal{F}_{p}$ and train only the classification module $\mathcal{F}_{c}$ for prediction, denoted as w/o Proj. To test the necessity of using LoRA Hu et al. (2022) for fine-tuning, we keep the rest of InterCLIP-MEP unchanged and freeze all self-attention weights of InterCLIP, denoted as w/o LoRA, always selecting the optimal memory size $L$ for the MEP during inference. To evaluate the effectiveness of the MEP, we train both $\mathcal{F}_{p}$ and $\mathcal{F}_{c}$ but use only $\mathcal{F}_{c}$ during inference, denoted as w/o MEP.
@@
Table 4 reports all results. All variants show performance declines compared to the baseline, demonstrating the importance of each module in InterCLIP-MEP. For w/ TW and w/o Inter, the w/o MEP variant performed worse than the w/o Proj variant. However, for w/ T2V and w/ V2T, the w/o MEP variant performs better than the w/o Proj variant. This suggests that backbones with strong image-text interaction capabilities benefit from training the classification module $\mathcal{F}_{c}$ along with the projection module $\mathcal{F}_{p}$, even without using MEP during inference. We also find that not using LoRA to fine-tune the self-attention modules results in significant performance loss, indicating that the original CLIP’s vanilla space is not directly suitable for the sarcasm detection task.
Case study.